# Segment Anything Model with Uncertainty Rectification for Auto-Prompting Medical Image Segmentation

### 13 Dec 2023 v2

@@ 3.4 Uncertainty Rectification Module
where $S_{\hat{y}}$ and $S_{b}$ represent the area of the initial segmentation mask and corresponding bounding box, respectively. Lower ratio indicates that the target occupies a smaller portion of the bounding box and there might be more uncertainty in regions surrounding the target including potential false positives or false negatives in the segmentation result. Then the high-uncertainty areas are selected.
@@
1:  Given an input image $x$
2:  Locate extreme points of the target organ and generate bounding box prompt $b$.
3:  Augmentation of initial prompt $b$ to generate $B$ consists of $\{b^{1},b^{2},\dots,b^{n}\}$ based on pre-defined number and ratio.
4:  Generate segmentation outputs $\{y^{1},y^{2},\dots,y^{n}\}$ based on multi-prompt inputs.
5:  Generate ensembled segmentation result $\hat{y}$ and estimated uncertainty $u(\hat{y})$ as 2 and 3
6:  Generate high uncertainty mask $M_{unc}$ with class-specific threshold as 4 and 5.
@@
14:  return  Final rectified segmentation result $\hat{y}_{r}$
By confidence-based filtering, we can divide the image into three parts, $M_{t}$ represents the certain regions of target inside high uncertainty areas, $M_{b}$ represents the certain regions of background outside the uncertainty areas, and $M_{unc}$ represents the uncertainty areas with high uncertainty. For rectification of uncertain regions, we estimate the average intensity of the image within certain regions of target and background as $I_{t}$ and $I_{b}$. If the pixel intensity of $M_{unc}$ is within a certain range of image intensity included in $M_{t}$, it is included to be part of the final segmentation result. The overall workflow is illustrated in 1.
Table 2: Ablation analysis of different components of uncertainty rectification approach.
Dataset StructSeg FLARE 22
@@
Table 3: Quantitative evaluation of Dice Similarity Coefficient (DSC) of different rectification methods with comparison to auto-prompting and manual prompting settings using SAM / MedSAM backbone for 3D head-and-neck organ segmentation in the StructSeg dataset. Higher values represent better segmentation performance.
@@
Table 4: Quantitative evaluation of Dice Similarity Coefficient (DSC) of different rectification methods with comparison to auto-prompting and manual prompting settings using SAM / MedSAM backbone for 3D abdominal organ segmentation in the FLARE 22 dataset. Higher values represent better segmentation performance.
@@
4 Experiments
4.1 Dataset and Model
We conduct experiments on two different medical image segmentation datasets. The first dataset is Automatic Structure Segmentation for Radiotherapy Planning Challenge Task1 dataset (StructSeg), which contains 50 CT scans for the segmentation of 22 head-and-neck (HaN) organs including brain stem, left eye, right eye, left lens, right lens, left optic nerve, right optic nerve, optic chiasma, left temporal lobes, right temporal lobes, pituitary, left parotid gland, right parotid gland, left inner ear, right inner ear, left middle ear, right middle ear, left TM joint, right TM joint, spinal cord, left mandible and right mandible. The second dataset is Fast and Low-resource semi-supervised Abdominal oRgan sEgmentation Challenge labeled set (FLARE 22), which contains 50 CT scans for the segmentation of 13 abdominal organs including liver, right kidney, spleen, pancreas, aorta, inferior vena cava, right adrenal gland, left adrenal gland, gallbladder, esophagus, stomach, duodenum, left kidney. These two datasets are highly representative as they contain the vast majority of important organs in the human body, providing good coverage for organ structure segmentation tasks in common clinical scenarios. To validate the effectiveness of our proposed framework, we integrate two different segmentation models with ViT-B backbone for the experiments: the original SAM [16] and MedSAM [23], a specialized variant of SAM fine-tuned on medical image datasets.
4.2 Implementation Details
Following the settings in [18], we randomly select five scans as support images for both datasets. For each organ in these scans, we compute the maximum and minimum coordinates and take the average of the coordinates and features across the support images to generate an average representation of latent coordinates and feature extreme points for auto-prompting. Specifically, the voxel spacing of input images is re-scaled to $3\times 3\times 3$ $mm^{3}$ with cropping patch sizes of $64\times 64\times 64$ pixels. We extend the 3D bounding box obtained from localization by [2, 10, 10] pixels in the z, x, and y directions, respectively to ensure that the targeted organ was completely encapsulated within the box. For data preprocessing, we perform the same procedure as described in [23], which includes adjusting the slice resolution to 3 × 1024 × 1024 and normalizing all images to the range [-500, 1000] as this range effectively encompasses most tissues. In addition to the automatic localization to generate prompts, we also conduct a series of experiments for manual prompting-based interactive segmentation, which are simulated based on the ground truth. To mimic the potential inaccuracies in manually drawn bounding boxes, we introduce a random perturbation of 0-20 pixels. In our experiments, we use the Dice Similarity Coefficient (DSC) as the evaluation metric of the segmentation task.
$DSC(G,S)=\frac{2|G\cap S|}{|G|+|S|}$ (6)
@@ 4.2 Implementation Details
4.3 Ablation Analysis
In this section, we aim to evaluate the effectiveness of our uncertainty rectification approach. As shown in Table 1, we first conduct experiments using different augmentation numbers and perturb ratios for prompt augmentation to select the optimal parameters for uncertainty estimation. From the results, we observe that incorporating perturbations into input prompts can improve the segmentation performance to some extent and enhance the model’s overall robustness. However, the benefits of additional augmentations were not observed when the number of augmentations exceeded three, or in some cases, even led to decreased segmentation performance. It is noteworthy that high perturbation ratios may cause some parts of the target to fall outside of the bounding box, resulting in decreased segmentation performance. Therefore, the augmentation ratio should be set in a suitable range. Table  1 presents the ablation analysis of different components of our framework. It can be seen that using class-specific thresholds for rectification increases the segmentation performance in both datasets.
4.4 Comparison Experiments
@@
We first conduct experiments on auto-prompting setting where the landmark detection model introduced in 3.1 is utilized to detect extreme points and generate bounding boxes as prompts for segmentation. We compare our rectification method with the FN/FP correction strategy in [37], where high-uncertainty areas are assumed to be potential false negative and positive regions for correction by removing potential false positive regions (Unc-FPC), adding potential false negative regions (Unc-FNC) and correcting both potential false negative and positive regions (Unc-FPNC) from the initial segmentation result. In addition to the auto-prompting, we also include manual prompting simulated based on the ground-truth masks for comparison. The experimental results of 3D head-and-neck organ segmentation in shown in Table 3. The head-and-neck region typically comprises relatively small organs with irregular shapes. We observe that directly applying SAM fails to segment target organs like the parotid gland (PG-L/PG-R) and spinal cord (SC), with an average dice coefficient lower than 10%, while ensembled results with augmented prompts can significantly improve the performance of these targets. For these small organs, the precise location of region of interest is crucial for the accuracy of subsequent segmentation and prompt augmentation can somehow ease the negative influence of generated inaccurate prompts. For 3D abdominal organ segmentation in Table 4, most target organs can be successfully segmented when directly applying SAM, especially for organs with neat shapes and well-defined boundaries like the liver and kidneys. However, we observe that prompt augmentation cannot consistently enhance performance and may even decrease the performance for some targets, since part of the organ may extend beyond the augmented bounding box in some instances. For rectification results, we observe that the performance of comparing strategies varied for different target organs due to the complex shape differences. Given that regions with high uncertainty may not necessarily correspond to false positives or negatives, especially for complex organs with differing shapes and volumes, this approach did not consistently improve performance. In comparison, our rectification strategy demonstrates more robust improvements for most segmentation targets, resulting in overall superior performance. We compare the performance of original and uncertainty rectified segmentation results in Fig 2 and 3. It can be observed that our framework further improves the segmentation performance with up to 10.7 % for head-and-neck organ segmentation and 13.8 % for abdominal organ segmentation in average dice coefficient. Fig. 4 presents the visual analysis of our proposed framework. The four columns represent the ground truth, ensembled segmentation result, high uncertainty mask, and rectified segmentation result, respectively. We can observe that rectified segmentation results have superior performance with less segmentation error compared with ensembled segmentation results, demonstrating the effectiveness of our proposed framework.
5 Discussion and Conclusion
@@
In this work, we present UR-SAM, an uncertainty rectified SAM framework for auto-prompting medical image segmentation by leveraging prompt augmentation of generated bounding box prompts for uncertainty evaluation and utilizing estimated uncertainty for rectification of segmentation results to improve the accuracy. Furthermore, the uncertainty map can help identify potential segmentation errors and support further analysis, offering valuable guidance in areas where manual focus and refinement are required for clinicians. The framework could also be integrated into interactive mechanisms, allowing users to select and modify specific regions that require refinement. Besides, we observe that MedSAM underperforms SAM to some extent, which is in line with the observations in [18]. This indicates that despite being fine-tuned on relatively large-scale medical datasets, the model does not consistently surpass classic SAM when applied to previously unseen datasets. While our method has demonstrated significant improvements, there are aspects where further enhancements could be made. One notable limitation is that we use a relatively simple strategy to classify pixels in regions of high uncertainty, which relied solely on the intensity of the pixel and did not consider information from neighboring pixels. Consequently, it still yields unsatisfactory results when it comes to segmenting some challenging classes. Future work will focus on incorporating contrast learning of pixels between different classes and integrating more recent medical adaptations of SAM [3, 33] for more comprehensive evaluations. Besides, it would be interesting to utilize uncertainty in the training procedure to select out more representative regions to further enhance model performance [34, 31].

### 18 Mar 2024 v3

@@ III-D Uncertainty Rectification Module
where $S_{\hat{y}}$ and $S_{b}$ represent the area of the ensembled segmentation mask and corresponding bounding box, respectively. Lower ratio indicates that the target occupies a smaller portion of the bounding box and there might be more uncertainty in regions surrounding the target including potential false positives or false negatives in the segmentation result. Then the high-uncertainty areas are selected.
@@
1:  Input an image $x$
2:  Locate extreme points of the target organ and generate bounding box prompt $b$ and corresponding bounding box $S_{b}$.
3:  Augment the prompt $b$ to generate $B$ consists of $\{b^{1},b^{2},\dots,b^{n}\}$ based on pre-defined number and ratio.
4:  Generate segmentation outputs $\{y^{1},y^{2},\dots,y^{n}\}$ based on multi-prompt inputs $B$.
5:  Generate ensembled segmentation result $\hat{y}$, ensembled segmentation mask $S_{\hat{y}}$ and estimated uncertainty $u(\hat{y})$ as Eq.(1) and Eq.(2).
6:  Generate high uncertainty mask $M_{unc}$ with class-specific threshold as Eq.(3) and Eq.(4).
@@
14:  return  Rectified segmentation result $\hat{y}_{r}$
By confidence-based filtering, we can divide the image into three parts, $M_{t}$ represents the certain regions of target inside high uncertainty areas, $M_{b}$ represents the certain regions of background outside the uncertainty areas, and $M_{unc}$ represents the uncertain regions with high uncertainty. For rectification of uncertain regions, we estimate the average intensity of the image within certain regions of target and background as $I_{t}$ and $I_{b}$. If the pixel intensity of $M_{unc}$ is within a certain range of image intensity included in $M_{t}$, it is included to be part of the final segmentation result. The overall workflow is illustrated in 1.
IV Experiments
IV-A Dataset and Model
We conduct experiments on two different medical image segmentation datasets. The first dataset is Automatic Structure Segmentation for Radiotherapy Planning Challenge Task1 dataset (StructSeg) [^1] https://structseg2019.grand-challenge.org, which contains 50 CT scans for the segmentation of 22 head-and-neck (HaN) organs including brain stem (BS), left eye (E-L), right eye (E-R), left lens (L-L), right lens (R-L), left optic nerve (ON-L), right optic nerve (ON-R), optic chiasma (OC), left temporal lobes (TL-L), right temporal lobes (TL-R), pituitary (P), left parotid gland (PG-L), right parotid gland (PG-R), left inner ear (IE-L), right inner ear (IE-R), left middle ear (ME-L), right middle ear (ME-R), left TM joint (J-L), right TM joint (J-R), spinal cord (SC), left mandible (M-L) and right mandible(M-R). The second dataset is the labeled set of Fast and Low-resource semi-supervised Abdominal oRgan sEgmentation Challenge (FLARE 22) [47], which contains 50 CT scans for the segmentation of 13 abdominal organs including liver, right kidney, spleen, pancreas, aorta, inferior vena cava (IVC), right adrenal gland (RAG), left adrenal gland (LAG), gallbladder (Gall), esophagus, stomach, duodenum, and left kidney. These two datasets are highly representative as they contain the vast majority of important organs in the human body, providing good coverage for organ structure segmentation tasks in common clinical scenarios. To validate the effectiveness of our proposed framework, we integrate two different segmentation models with ViT-B backbone for the experiments: the original SAM [9] and MedSAM [19], a specialized variant of SAM fine-tuned on medical image datasets.
IV-B Implementation Details and Evaluation Metrics
All of our experiments are conducted using an NVIDIA A100 GPU. Following the settings in [16], we randomly select five scans as support images for both datasets. For each organ in these scans, we compute the maximum and minimum coordinates and take the average of the coordinates and features across the support images to generate an average representation of latent coordinates and feature extreme points for auto-prompting. Specifically, the voxel spacing of input images is re-scaled to $3\times 3\times 3$ $mm^{3}$ with cropping patch sizes of $64\times 64\times 64$ pixels. The 3D bounding box obtained from localization is extended by [2, 10, 10] pixels in the z, x, and y directions, respectively to ensure that the targeted organ is completely encapsulated within the box. For data preprocessing, we perform the same procedure as described in [19], which includes adjusting the slice resolution to 3 × 1024 × 1024 and normalizing all images to the range [-500, 1000] as this range effectively encompasses most tissues. In addition to the automatic localization to generate prompts, we also conduct a series of experiments for manual prompting-based interactive segmentation, which are simulated based on the ground truth. To mimic the potential inaccuracies in manually drawn bounding boxes, we introduce a random perturbation of 0-20 pixels. In our experiments, we use the Dice Similarity Coefficient (DSC) as the evaluation metric of the segmentation task, which is one of the most commonly used evaluation metric for image segmentation to measure the degree of pixel-wise overlap between the segmentation results and the ground truth. Higher DSC indicate better segmentation performance. The calculation is defined as follows:
$DSC(G,S)=\frac{2|G\cap S|}{|G|+|S|}$ (5)
TABLE II: Ablation analysis of different components of uncertainty rectification approach.
Dataset StructSeg DSC FLARE22 DSC
@@ IV-B Implementation Details and Evaluation Metrics
TABLE III: Quantitative evaluation of Dice Similarity Coefficient (DSC) of different rectification methods with comparison to auto-prompting and manual prompting settings using SAM / MedSAM backbone for 3D head-and-neck organ segmentation in the StructSeg dataset. Higher values represent better segmentation performance.
@@
IV-C Ablation Analysis
In this section, we aim to evaluate the effectiveness of our uncertainty rectification approach. As shown in Table I, we first conduct experiments using different augmentation numbers and perturb ratios for prompt augmentation to select the optimal parameters for uncertainty estimation. From the results, we observe that incorporating perturbations into input prompts can improve the segmentation performance to some extent and enhance the model’s overall robustness. However, the benefits of additional augmentations were not observed when the number of augmentations exceeded three, or in some cases, even led to decreased segmentation performance. It is noteworthy that high perturbation ratios may cause some parts of the target to fall outside of the bounding box, resulting in decreased segmentation performance. Therefore, the augmentation ratio should be set in a suitable range. Table II presents the ablation analysis of different components of our framework. It can be seen that using class-specific thresholds for rectification increases the segmentation performance in both datasets.
TABLE IV: Quantitative evaluation of Dice Similarity Coefficient (DSC) of different rectification methods with comparison to auto-prompting and manual prompting settings using SAM / MedSAM backbone for 3D abdominal organ segmentation in the FLARE 22 dataset. Higher values represent better segmentation performance.
@@ IV-C Ablation Analysis
TABLE V: Comparison of our method with state-of-the-are specialist segmentation models under the same experimental settings. Specifically, N denotes the count of slices containing the target object.
Methods Manual Annotation Manual Prompting StructSeg DSC
nnUNet-2D [8] 5 labeled images - 0.422
nnUNet-3D [8] 5 labeled images - 0.557
SAM [9] - N bboxes for each target 0.489
MedSAM [19] - N bboxes for each target 0.553
MedLSAM (SAM backbone) [16] 5 support images - 0.394
MedLSAM (MedSAM backbone) [16] 5 support images - 0.453
UR-SAM (SAM backbone) 5 support images - 0.501
UR-SAM (MedSAM backbone) 5 support images - 0.486
IV-D Comparison Experiments
We first conduct experiments on auto-prompting setting where the landmark detection model introduced in Sec. III-A is utilized to detect extreme points and generate bounding boxes as prompts for segmentation. We compare our rectification method with the FN/FP correction strategy in [46], where high-uncertainty areas are assumed to be potential false negative and positive regions for correction by removing potential false positive regions (Unc-FPC), adding potential false negative regions (Unc-FNC) and correcting both potential false negative and positive regions (Unc-FPNC) from the initial segmentation result. In addition to the auto-prompting, we also include manual prompting simulated based on the ground-truth masks for comparison.
@@ IV-D Comparison Experiments
The experimental results of 3D head-and-neck organ segmentation in shown in Table III. The head-and-neck region typically comprises relatively small organs with irregular shapes. We observe that directly applying SAM fails to segment target organs like the spinal cord (SC) and parotid gland (PG-L / PG-R), with an average dice coefficient lower than 10%, while ensembled results with augmented prompts can significantly improve the performance of these targets. For these small organs, the precise location of region of interest is crucial for the accuracy of subsequent segmentation and prompt augmentation can somehow ease the negative influence of generated inaccurate prompts. For 3D abdominal organ segmentation in Table IV, most target organs can be successfully segmented when directly applying SAM, especially for organs with neat shapes and well-defined boundaries like the liver and kidneys. However, we observe that prompt augmentation cannot consistently enhance performance and may even decrease the performance for some targets, since part of the organ may extend beyond the augmented bounding box in some instances.
For rectification results, we observe that the performance of comparing strategies varied for different target organs due to the complex shape differences. Given that regions with high uncertainty may not necessarily correspond to false positives or negatives, especially for complex organs with differing shapes and volumes, this approach did not consistently improve performance. In comparison, our rectification strategy demonstrates more robust improvements for most segmentation targets, resulting in overall superior performance. We compare the performance of original and uncertainty rectified segmentation results in Fig. 3 and 4. It can be observed that our framework further improves the segmentation performance with up to 10.7 % for head-and-neck organ segmentation and 13.8 % for abdominal organ segmentation in average dice coefficient. Fig. 5 presents the visual analysis of our proposed framework. The four columns represent the ground truth, ensembled segmentation result, high uncertainty mask, and rectified segmentation result, respectively. We can observe that rectified segmentation results have superior performance with less segmentation error compared with ensembled segmentation results, demonstrating the effectiveness of our proposed framework.
In Table V, we make a comparative analysis of SAM-based auto-prompting segmentation methods and interactive segmentation methods based on manual prompting with training-based state-of-the-art automatic segmentation methods like nnUNet [8]. Despite achieving zero-shot segmentation without the need for voxel-wise annotations, both the original SAM [9] and its medical adaptions [19] require manual prompting of each target organ for each testing case to obtain the segmentation, which directly increase the burden for applications. Compared with training-based task-specific segmentation methods, SAM-based auto-prompting methods can leverage the knowledge base of foundation models for automatic segmentation, requiring training only a lightweight localization model. For instance, the training of 3D nnUNet needs around $\sim$ 25 hours, while the localization model only $\sim$ 30 minutes.
V Discussion and Conclusion
@@
In this work, we present UR-SAM, an uncertainty rectified SAM framework for auto-prompting medical image segmentation by leveraging prompt augmentation of generated bounding box prompts for uncertainty evaluation and utilizing estimated uncertainty for rectification of segmentation results to improve the accuracy. Furthermore, the uncertainty map can help identify potential segmentation errors and support further analysis, offering valuable guidance in areas where manual focus and refinement are required for clinicians. The framework could also be integrated into interactive mechanisms, allowing users to select and modify specific regions that require refinement. Besides, we observe that MedSAM underperforms SAM to some extent, which is in line with the observations in [16]. This indicates that despite being fine-tuned on relatively large-scale medical datasets, the model does not consistently surpass classic SAM when applied to previously unseen datasets.
While our method has demonstrated significant improvements, there are aspects where further enhancements could be made. One notable limitation is that we use a relatively simple strategy to classify pixels in regions with high uncertainty, which relied solely on the intensity of the pixel and did not consider information from neighboring pixels. Consequently, it still yields unsatisfactory results when it comes to segmenting some challenging classes. Future work will focus on incorporating contrast learning of pixels between different classes for more comprehensive evaluations. Besides, it would be interesting to utilize uncertainty in the training procedure to select out more representative regions to further enhance model performance [48, 49].